OpenStudy (anonymous):

what is the sum of the first 41 even numbers

Vocaloid (vocaloid):

use the sum formula for arithmetic sequences S = (n/2)(2a + (n-1)d) n = 41, a = 0, d = 2

OpenStudy (loser66):

\(\sum_{n=1}^{41} 2n = 2\sum_{n=1}^{41} n = 2*\dfrac{41(41+1)}{2}= 41*42 =....\)